OpenStudy (anonymous):

eleventh root of x to the tenth power over eleventh root of x to the sixth power??

OpenStudy (anonymous):

\[\frac{ \sqrt[11]{x}^{10} }{ \sqrt[11]{x}^{6} } = \frac{ ({x}^{1/11})^{10} }{ ({x}^{1/11})^{6} }=({x}^{1/11})^{4}={x}^{4/11}\]Let me know if you would like a more detailed explanation.